# Near instantaneous O(1) Analog Solver Circuit for Linear Symmetric Positive-Definite Systems

### v1

@@ Abstract
Accelerating the solution of linear systems of equations is critical due to their central role in numerous applications, such as scientific simulations, data analytics, and machine learning. This paper presents a general-purpose analog direct solver circuit designed to accelerate the solution of positive definite symmetric linear systems of equations. The proposed design leverages non-inverting operational amplifier configurations to create a negative resistance circuit, effectively modeling any symmetric system. The paper details the principles behind the design, optimizations of the system architecture, and numerical results that demonstrate the robustness of the design. The findings reveal that the proposed system solves diagonally dominant symmetric matrices with $O(1)$ complexity, achieving the theoretical maximum speed as the circuit relies solely on resistors. For non-diagonally dominant symmetric positive-definite systems, the solution speed depends on matrix properties such as eigenvalues and the maximum off-diagonal term, but remains independent of matrix size.
Index Terms:
analog computing, linear algebra, resistive networks
@@ I Introduction
Solving systems of linear equations are fundamental for addressing real-world challenges across various domains, especially in the era of big data where the volume and complexity of datasets continue to grow exponentially. Although powerful, traditional digital computing approaches are encountering limitations in parallelization and scalability, particularly as Moore’s Law is close to reaching its asymptotic plateau. Thus, accelerating the solution of linear systems of equations can be immensely impactful for applications where the ability to process vast amounts of data efficiently is paramount including scientific simulations, data analytics, machine learning and several others.
@@ I-A Existing Paradigms for Solving Large Systems of Equations
For decades, the primary engine for tackling these problems has been digital computing. Classical direct factorization methods, such as Gaussian elimination, LU decomposition, and Cholesky factorization, offer exact solutions but typically exhibit a computational complexity of $O(n^{3})$, where $n$ is the number of unknowns [1]. Iterative methods, including Jacobi, Gauss-Seidel, Successive Over-Relaxation (SOR), and Conjugate Gradient (CG), offer alternatives, particularly for large, sparse systems [2]. While often exhibiting lower complexity $O(n^{2})$ per iteration, their convergence rate depends heavily on matrix properties such as condition number [2]. As system size ($n$) explodes in modern applications, digital solvers, both direct and iterative, face significant bottlenecks. Executed on traditional von Neumann architectures, they have substantial time and energy overheads associated with data movement between memory and processing units, pushing power consumption and execution time to prohibitive levels [3]. This challenge motivates the exploration of fundamentally different computing paradigms such as quantum computing and analog computing.
@@
With quantum computing, algorithms like the Harrow-Hassidim-Lloyd (HHL) algorithm theoretically promise an exponential speedup for solving certain linear systems, potentially achieving complexity of $\text{log}(n)$ under specific conditions on the system matrix $A$ and the desired output precision[4]. Although its long-term promise holds, quantum computing faces challenges such as qubit coherence times and error rates. Thus, it is currently not a viable option for large-scale problems [5, 6].
@@
An alternative gaining significant attention is analog computing. Advancements in materials science and device physics have led to the development of analog Compute-In-Memory (CIM) architectures [7]. This paradigm shifts computation directly onto memory elements, drastically reducing data movement[8]. Advancements of non-volatile resistive memory technologies, particularly memristors (or Resistive RAM - ReRAM), organized into crossbar arrays led to the evolution of this architecture [9]. Memristors are variable resistor devices that can act as memory by changing their conductance values. Through ongoing advancements, the number of unique conductance values that a single memristor device can represent has progressively increased, with milestones demonstrating 2 bits [10], 5 bits [11], 9 bits [12], and ultimately 11 bits [13]. The precision of these unique values is limited by reading noise, writing variabilities, conductance drift, retention failure, and programming inaccuracies [14, 15, 16, 17, 18, 19]. Advances in techniques to overcome these challenges allowed a progressive increase in precision. Use of an error compensation technique across multiple (no more than five) arrays achieved double precision [20]. These advancements highlight the potential of analog computing to perform computations with accuracy comparable to digital computing.
@@
Analog computing systems, particularly those leveraging memristor-based compute-in-memory architectures, have shown promise in performing matrix-vector multiplication (MVM) efficiently, potentially reducing computational complexity and energy consumption [21]. The crossbar arrays in the architecture naturally implement MVM, a core operation in many linear system solvers, by leveraging fundamental physical principles like Ohm’s Law and Kirchhoff’s current law [22]. The complexity of MVM operation using memristors in the crossbar architecture is $O(1)$ compared to $O(n^{2})$ in digital computing [7]. Researchers report that a MVM CIM circuit can achieve 100x more energy efficiency than GPUs [23].
@@
Solving linear systems ($Ax=b$) using analog CIM has consequently gained a lot of attention, precisely because it relies heavily on MVM and demands high precision. These analog designs can be grouped into three classes: hybrid solvers, dynamical solvers, and direct feedback solvers. The hybrid (digital and analog) systems solve linear equations iteratively using MVM circuits. Progression of research in hybrid systems can be observed in the increasing integration of analog components within digital computers: ranging from the use of analog CIM as a quick generator of an initial guess for a digital solver [24], to the implementation of analog preconditioners [25], and to accelerating all MVM operations within conventional iterative algorithms like the generalized minimum residual (GMRES) [26]. These systems show potential in accelerating the MVM part of algorithms by reducing run-time by up to 1500x and saving energy by 8.5x compared to digital computing [24].
@@
The second class of analog designs is dynamical systems, which convert the system of linear equations to a system of differential equations (DEs) whose steady-state solution is the solution of the linear system of equations. Analog systems are transient systems best described using partial differential equations (PDEs) and ordinary differential equations (ODEs). Researchers use analog integrator circuits to solve the equivalent system of DEs [27, 28], while other researchers utilize transconductance devices in their designs to solve the DEs [29, 30]. The main drawback of these systems is that, by converting the linear system into DEs, the solution can have a long settling time as it depends on dynamic components such as operational amplifiers and capacitors.
@@
The third class of analog designs is the direct feedback methods. The goal of these systems is to achieve one-step solutions where the solution time is independent of the size of the system. Utilizing the MVM crossbar architecture in feedback using Transimpedance Amplifiers (TIA) can solve positive definite (PD) matrices in one step [31]. For sparse systems, the direct CIM design is shown to have a complexity that is independent of the matrix size ($n$) and given by $O(1/\lambda_{min})$ where $\lambda_{min}$ denotes the minimum eigenvalue of the matrix[32]. In comparison, the complexity of solving sparse matrices by digital computing using Conjugate Gradient method is $O(n)$ and by quantum computing it is $O(\kappa^{2}\text{log}(n))$ where $\kappa$ denotes the condition number of the matrix [32]. For more general matrices, complexity of the direct feedback analog computing is independent of matrix size but quadratic in the condition number ($O(\kappa^{2})$) [33]. These designs contain active components that can still result in long settling time that is problem-dependent.
@@
A common characteristic across many of these analog solver architectures, particularly those relying on dynamical systems or feedback loops incorporating active components, is the presence of inherent settling times. Whether due to explicit time constants (RC) in dynamical systems, the propagation delays, the finite gain-bandwidth product of operational amplifiers used in feedback and peripheral circuits, or the convergence time of the implemented algorithm itself, the analog computation requires long settling time to converge to the final solution. This settling time can depend on the system size, matrix properties (e.g., condition number), and circuit non-idealities, potentially limiting the achievable speedup, especially compared to the near-instantaneous potential of passive MVM.
@@ I-B Novel Approach Using Configurable Network of Resistors
Here we introduce a fundamentally different analog CIM design for solving symmetric linear systems of equations at unprecedented speed. Instead of modeling the mathematical process of solving a linear system of equations, the proposed analog design is constructed to directly map the given linear system of equations $Ax=b$ into a resistive network. The configuration and conductance values of this network are derived directly from the matrix $A$ and the vector $b$. The governing equations of the resistive network in this design are the given linear equations, and the set of nodal voltages in the network is the solution vector of the given linear system.
First, a preliminary circuit design based on direct mapping is developed to solve a $n\times n$ symmetric positive-definite (SPD) matrix system of equations. However, the speed and accuracy of the solution are limited by the number of positive off-diagonal values (up to $(n^{2}-n)/2$ values). The reason is that the circuit resulting from mapping the positive off-diagonal terms contains active components, namely operational amplifiers in a negative feedback loop, whereas the negative off-diagonal term is mapped directly into a passive resistor. The proposed design transforms the $n\times n$ SPD system to a $2n\times 2n$ SPD system with at most $n$ positive off-diagonal values without adding extra computations. Solving the transformed system is shown to require fewer components in the circuitry, speed up the convergence by orders of magnitude, consume less power, and allow the design to utilize the MVM architecture.
For the class of Symmetric Diagonally Dominant (SDD) systems, the mapping of the transformed system results in a purely passive resistive network, where the solution is achieved near-instantaneous $O(1)$. For the broader SPD class, the mapping of the transformed system results in a resistive network that includes active components, particularly operational amplifiers in negative feedback. The speed and accuracy of the solution in this case depend on the properties of the matrix $A$ such as the maximum off-diagonal term, but, critically, remain independent of the size of the matrix.
@@ II-A Equivalent Resistive Network
$\begin{bmatrix}k_{ij}&-k_{ij}\\ -k_{ij}&k_{ij}\end{bmatrix}\begin{Bmatrix}x_{i}\\ x_{j}\end{Bmatrix}=\begin{Bmatrix}i_{i}^{el}\\ i_{j}^{el}\end{Bmatrix},$ (3)
where $k_{ij}$ is the conductance of the resistor between nodes $i$ and $j$, $x_{i}$ and $x_{j}$ are the voltage at nodes $i$ and $j$, respectively, and $i_{i}^{el}$ and $i_{j}^{el}$ are the currents flowing out of nodes $i$ and $j$, respectively. Similarly, the current $i^{el}_{i}$ flowing out of node $i$ due to a resistor connected between node $i$ and the ground is obtained as follows:
$k_{0i}\ x_{i}=i_{i}^{el},$ (4)
where $k_{0i}$ is the conductance of the resistor between node $i$ and the ground. The resistive network satisfies Kirchhoff’s current law (KCL), requiring that the summation of the current flowing out of each node $i$ be equal to the external current $b_{i}$ flowing into the node. Assembling the resistor Eqs. (3)-(4), for each resistor in the network, one obtains the system in Eq. (2). This system is easily solved by simply measuring the voltage at all the nodes that represent the unknowns $x_{i}$.
Figure 1: Resistive network for a 3x3 system $Ax=b$. The unknown vector $x$ represents the voltage at the nodes, while the right hand side vector $b$ represents the external current going in to the nodes. The matrix $A$ represents the conductance of the resistors between the nodes.
In this formulation, the off-diagonal terms of the resulting matrix $A$ are the negative of the conductance values between the corresponding nodes, while the sum of each column of the matrix $A$ is the conductance value of the resistor connecting the corresponding node to the ground. Thus, given a symmetric system of equations $Ax=b$, the equivalent resistive network can be constructed by computing the conductance of the resistors as follows:
$k_{ij}$ $=-A_{ij}\>\>\>\>\>\>\>for\>\>i\neq j,$ (5)
$k_{0i}$ $=\sum_{j=1}^{n}A_{ji},$ (6)
where $A_{ij}$ is the element at row $i$ and column $j$ of the given matrix $A$. The representation of a given symmetric system of equations by a resistive network is shown in Figure 2. The above equations show that the conductance of the resistors can be either negative or positive. A positive conductance is represented with a standard resistor, while the negative conductance is represented by the negative resistance circuit shown in Section II-B.
Figure 2: Resistive network for a 3x3 system $Ax=b$. The unknown vector $x$ represents the voltage at the nodes, while the right hand side vector $b$ represents the external current going into the nodes. The matrix $A$ represents the conductance of the resistors between the nodes.
@@ II-B Negative Resistance Circuit
Negative resistance can be realized by designing a circuit that reverses the direction of current flow between two nodes. Such a concept has previously been employed to solve non-homogeneous differential equations [35]. To improve both accuracy and computational speed, we propose an alternative design. This proposed design extends the traditional Non-inverting operational amplifier configuration. The traditional configuration is used to reverse the current between a node and the ground, whereas the proposed design can be used to reverse the current between any two nodes.
@@
To create the two nodes $i^{\prime}$ and $j^{\prime}$, the Non-inverting Operational Amplifier configuration is used as shown in Figure 3 with negative feedback loop. The two resistors $R_{1}$ and $R_{2}$ are chosen to have equal conductance to have a gain of 2. The circuit in Figure 3 shows the amplifier circuits together with the two resistors $k_{ii}^{\prime}$ and $k_{jj}^{\prime}$. Buffer circuits are added to stop unaccounted currents from being drawn from/ supplied to the nodes where computed currents will be drawn from/ supplied to. The modified circuit that includes the two buffer circuits is shown in Figure 3.
Figure 3: Negative Resistance Circuit (a) without buffers $i_{ie}\neq 0,i_{je}\neq 0$ and (b) with buffers $i_{ie}=i_{je}=0$

### v2

# Analog Solver Circuit for Linear Symmetric Positive-Definite Systems at a Complexity Independent of Matrix Size

@@ Abstract
Accelerating the solution of linear systems of equations is critical due to their central role in numerous applications, such as numerical simulations, data analytics, and machine learning. This paper presents an analog solver circuit designed to accelerate the solution of symmetric positive definite (SPD) linear systems of equations. The proposed design leverages non-inverting operational amplifier configurations to create a negative resistance circuit, effectively modeling any symmetric system. The paper details the principles behind the design, optimizations of the system architecture, and numerical results that demonstrate the robustness of the design. The findings reveal that the proposed system solves symmetric diagonally dominant (SDD) matrices with $O(1)$ complexity, achieving the theoretical maximum speed as the circuit relies solely on resistors. For non-diagonally dominant SPD systems, the solution speed depends on matrix properties, specifically eigenvalues and diagonal dominance deviation, but remains independent of the size of the matrix.
@@ I Introduction
Solving systems of linear equations are fundamental for addressing real-world challenges across various domains, especially in this era of big data where the volume and complexity of datasets continue to grow exponentially. Although powerful, traditional digital computing approaches are encountering limitations in parallelization and scalability, particularly as Moore’s Law approaches its asymptotic plateau. Accelerating the solution of linear systems of equations will be immensely impactful for applications such as scientific simulations, data analytics, and machine learning.
@@ I-A Existing Paradigms for Solving Large Systems of Equations
For decades, the primary engine for solving systems of linear equations has been digital computing. Classical direct factorization methods, such as Gaussian elimination, LU decomposition, and Cholesky factorization, offer exact solutions but typically exhibit a computational complexity of $O(n^{3})$, where $n$ is the number of unknowns [9]. Iterative methods, including Jacobi, Gauss-Seidel, Successive Over-Relaxation (SOR), and Conjugate Gradient (CG), offer alternatives, particularly for large, sparse systems [36]. While often exhibiting lower complexity $O(n^{2})$ per iteration, the convergence rate of iterative methods depends heavily on matrix properties such as condition number [36]. As the size ($n$) of the systems are ever increasing in modern applications, digital solvers, both direct and iterative, face significant bottlenecks. Executed on traditional von Neumann architectures, they have costly time and energy overheads associated with data movement between memory and processing units, pushing power consumption and execution time to prohibitive levels [15]. This challenge motivates the exploration of fundamentally different computing paradigms such as quantum computing and analog computing.
@@
With quantum computing, algorithms like the Harrow-Hassidim-Lloyd (HHL) algorithm theoretically promise an exponential speedup for solving certain linear systems, potentially achieving solver complexity of $\text{log}(n)$ under specific conditions on the system matrix $A$ and the desired output precision[13]. In such approaches, obtaining explicit solution components involves a separate readout process with linear complexity $O(n)$. Although its long-term promise holds, quantum computing faces challenges such as qubit coherence times and error rates. Thus, it is currently not a viable option for large-scale problems [31, 12].
@@
An alternative gaining significant attention is analog computing. Advances in materials science and device physics have led to the development of analog Compute-In-Memory (CIM) architectures [37]. This paradigm shifts computation directly onto memory elements, drastically reducing data movement[18]. Advancements of non-volatile resistive memory technologies, particularly memristors (or Resistive RAM - ReRAM), organized into crossbar arrays led to the evolution of this architecture [39]. Memristors are variable resistor devices that can act as memory by changing their conductance values. Through ongoing research, the number of unique conductance values that a single memristor device can represent has progressively increased, with milestones demonstrating 2 bits [43], 5 bits [19], 9 bits [6], and ultimately 11 bits [33]. The precision of these unique values is limited by reading noise, writing variabilities, conductance drift, retention failure, and programming inaccuracies [7, 20, 1, 27, 3, 4]. Techniques to overcome these challenges allowed a progressive increase in precision. Use of an error compensation technique across multiple (no more than five) arrays achieved double precision [38]. These advancements highlight the potential of analog computing to perform computations with accuracy comparable to digital computing.
@@
Analog computing systems, particularly those leveraging memristor-based compute-in-memory architectures, have shown promise in performing matrix-vector multiplication (MVM) efficiently, potentially reducing computational complexity and energy consumption [22]. The crossbar arrays in the architecture naturally implement MVM, a core operation in many linear system solvers, by leveraging fundamental physical principles like Ohm’s Law and Kirchhoff’s current law [16]. The complexity of MVM operation using memristors in the crossbar architecture is $O(1)$ compared to $O(n^{2})$ in digital computing [37]. Researchers report that a MVM CIM circuit can achieve 100x better energy efficiency than GPUs [40].
@@
Solving linear systems ($Ax=b$) using analog CIM has consequently gained a lot of attention, precisely because it relies heavily on MVM and demands high precision. These analog designs can be grouped into three classes: hybrid solvers, dynamical solvers, and feedback solvers. The hybrid (digital + analog) systems solve linear equations iteratively using MVM circuits. Progress in hybrid systems is observed in the increasing integration of analog components within digital computers: ranging from the use of analog CIM as a quick generator of an initial guess for a digital solver [35], to the implementation of analog preconditioners [8], and to accelerating all MVM operations within conventional iterative algorithms such as the generalized minimum residual (GMRES) [21]. These systems show potential in accelerating the MVM part of algorithms by reducing run time by up to 1500x and saving energy by 8.5x compared to digital computing [35].
@@
The second class of analog designs is dynamical systems, which convert the system of linear equations to a system of differential equations (DEs) whose steady-state solution is the solution of the linear system of equations. Analog systems are transient systems best described using partial differential equations (PDEs) and ordinary differential equations (ODEs). Researchers use analog integrator circuits to solve the equivalent system of DEs [45, 17], while other researchers utilize transconductance devices in their designs to solve the DEs [14, 30]. To improve scalability, recent work has also proposed integrator circuits based on ring oscillators, which can reduce both area and power consumption compared to conventional operational amplifier (OA)-based integrators [23]. The main drawback of these systems is that, by converting the linear system into DEs, the solution can have a long settling time as it depends on dynamic components such as operational amplifiers and capacitors.
@@
The third class of analog designs is the feedback methods. These systems model iterative algorithms into feedback networks. The goal of these systems is to achieve one-step solutions where the solution time is independent of the size of the system. Using the MVM crossbar architecture in feedback with Operational Amplifiers (OAs) can solve positive definite (PD) matrices in one step [41]. For sparse systems, the direct CIM design is shown to have a complexity that is independent of the matrix size ($n$) and given by $O(1/\lambda_{min})$ where $\lambda_{min}$ denotes the minimum eigenvalue of the matrix[42]. In comparison, the complexity of solving sparse matrices by digital computing using Conjugate Gradient method is $O(n)$ and by quantum computing it is $O(\kappa^{2}\text{log}(n))$ where $\kappa$ denotes the condition number of the matrix [42]. For more general matrices, complexity of the feedback analog computing is independent of matrix size but quadratic in the condition number ($O(\kappa^{2})$) [29]. In addition, these systems accelerate computing the inverse of the system matrix at a complexity of $O(n^{2})$ compared to a complexity of $O(n^{3})$ for digital computing [24]. One notable advantage of these analog designs is that they allow the use of the same architecture for multiple matrix operations [28]. To achieve higher speed and reduce the physical space of the design, researchers aim to reduce the number of active components needed in the design [25]. However, these designs always contain active components that can still result in long settling time, which is problem-dependent.
@@
A common characteristic across many of these analog solver architectures, particularly those relying on dynamical systems or feedback loops incorporating active components, is the presence of inherent settling times. Whether due to explicit time constants (RC) in dynamical systems, the propagation delays, the finite gain-bandwidth product of OAs used in feedback and peripheral circuits, or the convergence time of the implemented algorithm itself, analog computation requires long settling time to converge to the final solution. This settling time can depend on the system size, matrix properties (such as condition number), and circuit non-idealities, potentially limiting the achievable speedup, especially compared with the near-instantaneous potential of passive MVM.
@@ I-B Novel Approach Using Configurable Network of Resistors
Symmetric positive definite (SPD) linear systems arise in a wide range of scientific and engineering computations. Prominent examples include finite element solvers for solid mechanics, optimization of convex problems, and covariance-driven computations in filtering and uncertainty quantification. Because these applications often require solving large SPD systems repeatedly, accelerating SPD solvers is critical.
Here, we introduce a fundamentally different analog CIM design to solve symmetric positive definite (SPD) linear systems of equations at unprecedented speed. Instead of modeling the mathematical process of solving a linear system of equations, the proposed analog design is constructed to directly map the given linear system of equations $Ax=b$ into a resistive network. The configuration and conductance values of this network are derived directly from the matrix $A$ and the vector $b$. The governing equations of the resistive network in this design are the given linear equations, and the set of nodal voltages in the network is the solution vector of the given linear system.
To illustrate, first, a preliminary circuit design based on direct mapping is developed to solve a $n\times n$ symmetric positive-definite (SPD) matrix system of equations. However, the speed and accuracy of the solution are limited by the number of active positive off-diagonal entries (up to $(n^{2}-n)/2$ entries). The reason is that the circuit resulting from mapping the positive off-diagonal terms contains active components, namely OAs in a negative feedback loop, whereas the negative off-diagonal terms are mapped directly into passive resistors. The proposed design transforms the $n\times n$ SPD system to a $2n\times 2n$ SPD system with at most $n$ positive off-diagonal values without adding extra computations. Solving the transformed system is shown to require fewer components in the circuitry, speed up the convergence by orders of magnitude, consume less power, and allow the design to utilize the MVM architecture.
For the class of Symmetric Diagonally Dominant (SDD) systems, the mapping of the transformed system results in a purely passive resistive network, where the solution is achieved at a near-instantaneous $O(1)$ rate. For the broader SPD class, the mapping of the transformed system results in a resistive network that includes active components, particularly OAs in negative feedback. The speed and accuracy of the solution in this case depend on the properties of the matrix $A$, such as the smallest eigenvalue, but, critically, remain independent of the size of the matrix.
@@ II-A Equivalent Resistive Network
Figure 1: A resistive network example for a 3x3 system $Ax=b$. The unknown vector $x$ represents the voltage at the nodes, while the right hand side vector $b$ represents the external current going in to the nodes. The matrix $A$ represents the conductance of the resistors between the nodes.
@@
$\begin{bmatrix}k_{ij}&-k_{ij}\\ -k_{ij}&k_{ij}\end{bmatrix}\begin{Bmatrix}x_{i}\\ x_{j}\end{Bmatrix}=\begin{Bmatrix}i_{i}^{el}\\ i_{j}^{el}\end{Bmatrix}$ (3)
where $k_{ij}$ is the conductance of the resistor between nodes $i$ and $j$, $x_{i}$ and $x_{j}$ are the voltage at nodes $i$ and $j$, respectively, and $i_{i}^{el}$ and $i_{j}^{el}$ are the current flowing out of nodes $i$ and $j$, respectively. Similarly, the current $i^{el}_{i}$ flowing out of node $i$ due to a resistor connected between node $i$ and the ground is obtained as follows:
$k_{0i}\ x_{i}=i_{i}^{el}$ (4)
where $k_{0i}$ is the conductance of the resistor between node $i$ and the ground. The resistive network satisfies Kirchhoff’s current law (KCL), requiring that the summation of the current flowing out of each node $i$ be equal to the external current $b_{i}$ flowing into the node. Assembling the resistor Equations (3)-(4), for each resistor in the network, one obtains the system in Equation (2). This system is easily solved by simply measuring the voltage at all the nodes that represent the unknowns $x_{i}$.
Figure 2: Theoretical construction of a resistive network for a given 3x3 system $Ax=b$. The unknown vector $x$ represents the voltage at the nodes, while the right hand side vector $b$ represents the external current going into the nodes. Conductance of the resistors are obtained from the components of $A$.
In this formulation, the off-diagonal terms of matrix $A$ are the negative of the conductance values between the corresponding nodes, and the sum of each column of the matrix $A$ is the conductance value of the resistor connecting the corresponding node to the ground, i.e.
$k_{ij}$ $=-A_{ij}\>\>\>\>\>\>\>for\>\>i\neq j$ (5)
$k_{0i}$ $=\sum_{j=1}^{n}A_{ji}$ (6)
where $A_{ij}$ is the element at row $i$ and column $j$ of the given matrix $A$. Thus, given a symmetric system of equations $Ax=b$, the equivalent resistive network can be constructed by computing the conductance of the resistors in circuit using Equations (5)-(6). For example, a given 3x3 symmetric system of equations $Ax=b$ can be represented by a resistive network as shown in Figure 2. Note that the conductance of the resistors resulting from Equations (5)-(6) can either be negative or positive. A positive conductance is represented with a standard resistor, while the negative conductance is represented by the negative resistance circuit shown in Section II-B.
@@ II-B Negative Resistance Circuit
Negative resistance can be realized by designing a circuit that reverses the direction of current flow between two nodes [5]. Here, we extend the traditional non-inverting operational amplifier configuration to reverse the current between any two nodes.
@@
To create the two nodes $i^{\prime}$ and $j^{\prime}$, the non-inverting operational amplifier configuration is used as shown in Figure 3 with negative feedback loop. The two resistors $R_{1}$ and $R_{2}$ are chosen to have equal conductance to have a gain of 2. The circuit in Figure 3 shows the amplifier circuits together with the two resistors $k_{ii}^{\prime}$ and $k_{jj}^{\prime}$. Buffer circuits are added to stop unaccounted currents from being drawn from / supplied to the nodes where computed currents will be drawn from / supplied to.
Figure 3: Negative Resistance Circuit with buffers $i_{ie}=i_{je}=0$
The governing differential equation of the proposed negative resistance design is derived from the transfer functions for the individual OAs. For the upper part of the circuit, the transfer functions in the $s$-domain for the buffer OA and the gain-2 OA are
$v_{i}(s)$ $=[x_{i}(s)-v_{i}(s)]L(s)$ (12)
$x_{j}^{\prime}(s)$ $=\left[x_{j}(s)-\frac{x_{j}^{\prime}(s)+v_{i}(s)}{2}\right]L(s)$ (13)
where $L(s)$ is the OA open-loop transfer function. The single-pole OA transfer function is defined as follows
$L(s)=\frac{L_{0}}{1+s/\omega_{0}}=\frac{1}{1/L_{0}+s/(L_{0}\omega_{0})}\approx\frac{L_{0}\omega_{0}}{s}$ (14)
where $L_{0}$ is the DC open-loop gain and $\omega_{0}$ is the 3-dB bandwidth [34]. Defining $\tau=1/(L_{0}\omega_{0})$ and combining Equations (12), (13) and (14) result in the following
$2\tau sx_{j}^{\prime}=2x_{j}-x_{j}^{\prime}-\frac{x_{i}}{1+\tau s}$ (15)
The equation is simplified to the following form in the $s$-domain
$(2\tau^{2}s^{2}+3\tau s+1)x_{j}^{\prime}=2x_{j}-x_{i}+2\tau sx_{j}$ (16)
Transforming the equation into the time domain gives the governing second-order differential equation as follows
$2\tau^{2}\frac{d^{2}x_{j}^{\prime}}{dt^{2}}+3\tau\frac{dx_{j}^{\prime}}{dt}+x_{j}^{\prime}=2x_{j}-x_{i}+2\tau\frac{dx_{j}}{dt}$ (17)
Another, further simplified differential equation that can be used to estimate the lower bound of the settling time is derived by assuming that the buffer OA is an ideal OA with infinite bandwidth. This assumption ignores the time delay in the buffer OA stage, where the transfer function in Equation (12) is modified to $v_{i}=x_{i}$. Following the same derivation steps, the resulting first-order governing differential equation for the lower bound of the settling time is as follows
$2\tau\frac{dx_{j}^{\prime}}{dt}+x_{j}^{\prime}=2x_{j}-x_{i}$ (18)